Problem 3 Let Xi, X2,... , Xn be a sequence of binary, i.i.d. random variables. Assume P (Xi 1) P (Xi = 0) = 1/2. Let Z be a parity check on seluence Xi, X2, ,X,, that is, Z = X BX2 e (a) Is Z statistically independent of Xi? (Assume n> 1) (b) Are X, X2, ..., Xn 1, Z statistically independent? (c) Are X, X2,.., Xn, Z statistically independent? (d) Is Z statistically independent of Xi if P (X,-1) #1/2? You lnay take n-2 here.
Answer #1

n- (a) P(7-1) Z = 1 X, takes 0 or 1, i = 1(1)n_ 1 but choice of Xn depends on X,-1 then P(Z = 0) = P(Xi = 1, Z = 1) = Similar

PlZ = 0) = P(Xi 1 , X2 = 0) + P(X1 = 0, X2 = 1) 2p(1-p) since Xis are iid So, 2 so, if pチ1/2, then PlXi=1.Z = 0) = p(1-p)メ2p(